# Kusatsu, Shiga 525-8577, Japan

### v1

@@ 3.2 Standard Courant sigma models with Courant algebroid gauging
$\vartheta$ $=z_{i}\delta x^{i}+p_{i}\delta q^{i}+\frac{1}{2}m_{ab}a^{a}\delta a^{b}$
$=z_{i}^{\nabla}\delta x^{i}+p_{i}\delta q^{i}+\frac{1}{2}m_{ab}a^{a}Da^{b},$ (38)
@@
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\bm{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+\frac{1}{2}m_{ab}{\bm{A}}^{a}D{\bm{A}}^{b}+{\bm{Z}}_{i}^{\nabla}{\bm{Q}}^{i}+\rho_{a}^{i}{\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}\right$.
$\left.+\frac{1}{3!}H_{ijk}(\bm{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{Q}}^{k}+\frac{1}{3!}T_{abc}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right],$ (52)
@@ 3.3 General Courant sigma models with Lie algebroid gauging
$\vartheta$ $=z_{i}\delta x^{i}+\frac{1}{2}k_{AB}\eta^{A}\delta\eta^{B}+w_{a}\delta a^{a}$
$=z_{i}^{\nabla}\delta x^{i}+\frac{1}{2}k_{AB}\eta^{A}\delta\eta^{B}+w_{a}Da^{a},$ (57)
@@
$S_{C}^{\nabla}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\bm{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+{\rho_{(E)}}^{i}_{A}{\bm{Z}}_{i}^{\nabla}{\bm{Y}}^{A}\right$.
$\left.\qquad+\frac{1}{3!}H_{ABC}(\bm{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{Y}}^{C}\right],$ (73)
@@
$S_{A}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[{\bm{W}}_{a}D{\bm{A}}^{a}+{\rho_{(A)}}^{i}_{a}(\bm{X}){\bm{Z}}_{i}{\bm{A}}^{a}+\frac{1}{2}C_{ab}^{c}{\bm{A}}^{a}{\bm{A}}^{b}{\bm{W}}_{c}\right]$
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[{\bm{W}}_{a}D{\bm{A}}^{a}+\rho^{i}_{a}(\bm{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}-\frac{1}{2}T_{ab}^{c}{\bm{A}}^{a}{\bm{A}}^{b}{\bm{W}}_{c}\right],$ (74)
where $D{\bm{A}}^{a}=\mathrm{d}{\bm{A}}^{a}-\omega^{a}_{bi}{\bm{A}}^{b}\mathrm{d}\bm{X}^{i}$. The total action functional is
@@
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\bm{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+{\bm{W}}_{a}D{\bm{A}}^{a}+{\rho_{(E)}}^{i}_{A}{\bm{Z}}_{i}^{\nabla}{\bm{Y}}^{A}\right$.
$\left.\quad+\frac{1}{3!}H_{ABC}(\bm{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{Y}}^{C}+{\rho_{(A)}}^{i}_{a}(\bm{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}-\frac{1}{2}T_{ab}^{c}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{W}}_{c}\right].$ (75)
@@ 3.4 General Courant sigma models with Courant algebroid gauging
where $Da^{a}=\delta a^{a}-\omega^{a}_{bi}a^{b}\delta x^{i}$. $\vartheta$ is invariant under changing of local coordinates on the graded manifold $T^{*}[2](E[1]\oplus A[1])$. Here transformations of local coordinates on a graded manifold $T^{*}[2]A[1]$ is given by Eqs. (213)–(215) for corresponding coordinates. Note that we do not consider covariant coordinate with respect to $T^{*}[2]E[1]$ by introducing a connection on $E$, since geometric structures of the sigma model, especially homological functions do not depend on the connection on $E$.
@@
Settings and results in Theorem 3.3 and 3.4 are geometric results connected to analysis of symmetries and reductions of Courant algebroids in [32, 33, 34, 35]. Especially, [35] uses graded geometry. We obtain QP-manifold and AKSZ sigma model realizations.
@@
$\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}\mathrm{d}\bm{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+\frac{1}{2}m_{ab}{\bm{A}}^{a}\mathrm{d}{\bm{A}}^{b}\right]$
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\bm{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+\frac{1}{2}m_{ab}{\bm{A}}^{a}D{\bm{A}}^{b}\right]$ (95)
@@
$S_{C}^{\nabla}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\bm{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+{\rho_{(E)}}^{i}_{A}{\bm{Z}}_{i}^{\nabla}{\bm{Y}}^{A}\right$.
$\left.\qquad+\frac{1}{3!}H_{ABC}(\bm{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{Y}}^{C}\right],$ (96)
@@
$S_{A}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[\frac{1}{2}m_{ab}(\bm{X}){\bm{A}}^{a}D{\bm{A}}^{b}+{\rho_{(A)}}^{i}_{a}(\bm{X}){\bm{Z}}_{i}{\bm{A}}^{a}+\frac{1}{3!}C_{abc}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right]$
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[\frac{1}{2}m_{ab}(\bm{X}){\bm{A}}^{a}D{\bm{A}}^{b}+{\rho_{(A)}}^{i}_{a}(\bm{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}+\frac{1}{3!}T_{abc}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right]$ (97)
where $D{\bm{A}}^{a}=\mathrm{d}{\bm{A}}^{a}-\omega^{a}_{bi}{\bm{A}}^{b}\mathrm{d}\bm{X}^{i}$. Therefore, the total action functional of the GCSM is
@@
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\bm{X}^{i}+\frac{1}{2}k_{AB}(\bm{X}){\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+\frac{1}{2}m_{ab}(\bm{X}){\bm{A}}^{a}D{\bm{A}}^{b}\right$.
$\left.+{\rho_{(E)}}^{i}_{A}{\bm{Z}}_{i}^{\nabla}{\bm{Y}}^{A}+\frac{1}{3!}H_{ABC}(\bm{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{Y}}^{C}+{\rho_{(A)}}^{i}_{a}(\bm{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}+\frac{1}{3!}T_{abc}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right].$ (98)
@@ 4.1 Introduction of fluxes
New Poisson brackets including $\Theta_{F}$} are as follows, Notation: $\partial_{[i}\alpha_{j]}=\partial_{i}\alpha_{j}-\partial_{j}\alpha_{i}$, and $\partial_{[i}T_{jk]}=\partial_{i}T_{jk}+\partial_{j}T_{ki}+\partial_{k}T_{ij}$. for an antisymmetric tensor $T_{ij}$.
${\{{{{\Theta_{A}}^{\nabla}},{{\Theta_{F}}}}\}}$ $=-\frac{1}{4}(\rho^{i}_{[a}\partial_{i}H^{(2)}_{ij|b]}-C_{ab}^{c}H^{(2)}_{ijc})q^{i}q^{j}a^{a}a^{b}+\frac{1}{3!}(\rho^{i}_{[a|}\partial_{i}H^{(1)}_{i|bc]}-C_{[ab}^{d}H^{(1)}_{i|d|c]})q^{i}a^{a}a^{b}a^{c}$
$\quad-\frac{1}{4!}(\rho^{i}_{[a}\partial_{i}H^{(0)}_{bcd}-C_{[ab}^{e}H^{(0)}_{e|cd]})a^{a}a^{b}a^{c}a^{d}$
$=-\frac{1}{4}(\rho^{k}_{a}\partial_{k}H^{(2)}_{ij|b]}-C_{ab}^{c}H^{(2)}_{ijc})q^{i}q^{j}a^{a}a^{b}+\frac{1}{3!}(\rho^{i}_{[a|}\partial_{k}H^{(1)}_{i|bc]}-C_{[ab}^{d}H^{(1)}_{i|d|c]})q^{i}a^{a}a^{b}a^{c}$
$\quad-\frac{1}{4!}(\rho^{i}_{[a}\partial_{k}H^{(0)}_{bcd}-C_{[ab}^{e}H^{(0)}_{e|cd]})a^{a}a^{b}a^{c}a^{d}.$ (101)
@@
${}^{A}\mathrm{d}^{\nabla}H^{(1)}-\mathrm{d}H^{(0)}=0.$ (113)
@@
$S_{F}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[\frac{1}{2}H^{(2)}_{ija}(\bm{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{A}}^{a}+\frac{1}{2}H^{(1)}_{iab}(\bm{X}){\bm{Q}}^{i}{\bm{A}}^{a}{\bm{A}}^{b}+\frac{1}{3!}H^{(0)}_{abc}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right].$ (115)
@@
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\bm{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+{\bm{W}}_{a}D{\bm{A}}^{a}+{\bm{Z}}_{i}^{\nabla}{\bm{Q}}^{i}+\rho^{i}_{a}(\bm{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}\right$.
$\left.+\frac{1}{3!}H^{(3)}_{ijk}(\bm{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{Q}}^{k}-\frac{1}{2}T_{ab}^{c}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{W}}_{c}\right$.
$\left.+\frac{1}{2}H^{(2)}_{ija}(\bm{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{A}}^{a}+\frac{1}{2}H^{(1)}_{iab}(\bm{X}){\bm{Q}}^{i}{\bm{A}}^{a}{\bm{A}}^{b}+\frac{1}{3!}H^{(0)}_{abc}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right].$ (116)

### v2

@@ 3.2 Standard Courant sigma models with Courant algebroid gauging
$\vartheta$ $=z_{i}\delta x^{i}-p_{i}\delta q^{i}-\frac{1}{2}m_{ab}a^{a}\delta a^{b}$
$=z_{i}^{\nabla}\delta x^{i}-p_{i}\delta q^{i}-\frac{1}{2}m_{ab}a^{a}Da^{b},$ (38)
@@
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\boldsymbol{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+\frac{1}{2}m_{ab}{\bm{A}}^{a}D{\bm{A}}^{b}+{\bm{Z}}_{i}^{\nabla}{\bm{Q}}^{i}+\rho_{a}^{i}{\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}\right$.
$\left.+\frac{1}{3!}H_{ijk}(\boldsymbol{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{Q}}^{k}+\frac{1}{3!}T_{abc}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right],$ (52)
@@ 3.3 General Courant sigma models with Lie algebroid gauging
$\vartheta$ $=z_{i}\delta x^{i}-\frac{1}{2}k_{AB}\eta^{A}\delta\eta^{B}-w_{a}\delta a^{a}$
$=z_{i}^{\nabla}\delta x^{i}-\frac{1}{2}k_{AB}\eta^{A}\delta\eta^{B}-w_{a}Da^{a},$ (57)
@@
$S_{C}^{\nabla}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\boldsymbol{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+{\rho_{(E)}}^{i}_{A}{\bm{Z}}_{i}^{\nabla}{\bm{Y}}^{A}\right$.
$\left.\qquad+\frac{1}{3!}H_{ABC}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{Y}}^{C}\right],$ (73)
@@
$S_{A}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[{\bm{W}}_{a}D{\bm{A}}^{a}+{\rho_{(A)}}^{i}_{a}(\boldsymbol{X}){\bm{Z}}_{i}{\bm{A}}^{a}+\frac{1}{2}C_{ab}^{c}{\bm{A}}^{a}{\bm{A}}^{b}{\bm{W}}_{c}\right]$
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[{\bm{W}}_{a}D{\bm{A}}^{a}+\rho^{i}_{a}(\boldsymbol{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}-\frac{1}{2}T_{ab}^{c}{\bm{A}}^{a}{\bm{A}}^{b}{\bm{W}}_{c}\right],$ (74)
where $D{\bm{A}}^{a}=\mathrm{d}{\bm{A}}^{a}-\omega^{a}_{bi}{\bm{A}}^{b}\mathrm{d}\boldsymbol{X}^{i}$. The total action functional is
@@
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\boldsymbol{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+{\bm{W}}_{a}D{\bm{A}}^{a}+{\rho_{(E)}}^{i}_{A}{\bm{Z}}_{i}^{\nabla}{\bm{Y}}^{A}\right$.
$\left.\quad+\frac{1}{3!}H_{ABC}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{Y}}^{C}+{\rho_{(A)}}^{i}_{a}(\boldsymbol{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}-\frac{1}{2}T_{ab}^{c}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{W}}_{c}\right].$ (75)
@@ 3.4 General Courant sigma models with Courant algebroid gauging
where $Da^{a}=\delta a^{a}-\omega^{a}_{bi}a^{b}\delta x^{i}$. $\vartheta$ is invariant under changing of local coordinates on the graded manifold $T^{*}[2](E[1]\oplus A[1])$. Here transformations of local coordinates on a graded manifold $T^{*}[2]A[1]$ is given by Eqs. (213)–(215) for corresponding coordinates. Note that we do not consider covariant coordinate with respect to $T^{*}[2]E[1]$ by introducing a connection on $E$, since geometric structures of the sigma model, especially homological functions do not depend on the connection on $E$. One can refer to Appendix (B.2) for curvatures, $E$-connections, $E$-torsions, basic curvatures, etc. in Courant algebroids.
@@
Settings and results in Theorem 3.3 and 3.4 are geometric results connected to analysis of symmetries and reductions of Courant algebroids in BursztynCavalcantiGualtieri ; BursztynCavalcantiGualtieri2 ; BursztynIglesiasPonteSevera ; Bursztyn:2023onv . Especially, Bursztyn:2023onv uses graded geometry. We obtain QP-manifold and AKSZ sigma model realizations.
@@
$\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}\mathrm{d}\boldsymbol{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+\frac{1}{2}m_{ab}{\bm{A}}^{a}\mathrm{d}{\bm{A}}^{b}\right]$
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\boldsymbol{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+\frac{1}{2}m_{ab}{\bm{A}}^{a}D{\bm{A}}^{b}\right]$ (95)
@@
$S_{C}^{\nabla}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\boldsymbol{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+{\rho_{(E)}}^{i}_{A}{\bm{Z}}_{i}^{\nabla}{\bm{Y}}^{A}\right$.
$\left.\qquad+\frac{1}{3!}H_{ABC}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{Y}}^{C}\right],$ (96)
@@
$S_{A}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[\frac{1}{2}m_{ab}(\boldsymbol{X}){\bm{A}}^{a}D{\bm{A}}^{b}+{\rho_{(A)}}^{i}_{a}(\boldsymbol{X}){\bm{Z}}_{i}{\bm{A}}^{a}+\frac{1}{3!}C_{abc}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right]$
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[\frac{1}{2}m_{ab}(\boldsymbol{X}){\bm{A}}^{a}D{\bm{A}}^{b}+{\rho_{(A)}}^{i}_{a}(\boldsymbol{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}+\frac{1}{3!}T_{abc}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right]$ (97)
where $D{\bm{A}}^{a}=\mathrm{d}{\bm{A}}^{a}-\omega^{a}_{bi}{\bm{A}}^{b}\mathrm{d}\boldsymbol{X}^{i}$. Therefore, the total action functional of the GCSM is
@@
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\boldsymbol{X}^{i}+\frac{1}{2}k_{AB}(\boldsymbol{X}){\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+\frac{1}{2}m_{ab}(\boldsymbol{X}){\bm{A}}^{a}D{\bm{A}}^{b}\right$.
$\left.+{\rho_{(E)}}^{i}_{A}{\bm{Z}}_{i}^{\nabla}{\bm{Y}}^{A}+\frac{1}{3!}H_{ABC}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{Y}}^{C}+{\rho_{(A)}}^{i}_{a}(\boldsymbol{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}+\frac{1}{3!}T_{abc}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right].$ (98)
@@ 4.1 Introduction of fluxes
New Poisson brackets including $\Theta_{F}$} are as follows,  [^2] Notation: $\partial_{[i}\alpha_{j]}=\partial_{i}\alpha_{j}-\partial_{j}\alpha_{i}$, and $\partial_{[i}T_{jk]}=\partial_{i}T_{jk}+\partial_{j}T_{ki}+\partial_{k}T_{ij}$ for an antisymmetric tensor $T_{ij}$, etc.
${\{{{{\Theta_{S}}^{\nabla}},{{\Theta_{F}}}}\}}$ $=-\frac{1}{3!}\nabla_{i}H^{(2)}_{jka}(x)q^{i}q^{j}q^{k}a^{a}+\frac{1}{4}\nabla_{i}H^{(1)}_{jab}(x)q^{i}q^{j}a^{a}a^{b}-\frac{1}{3!}\partial_{i}H^{(0)}_{abc}(x)q^{i}a^{a}a^{b}a^{c}.$ (100)
${\{{{{\Theta_{A}}^{\nabla}},{{\Theta_{F}}}}\}}$ $=-\frac{1}{4}(\rho^{i}_{a}\partial_{i}H^{(2)}_{ijb}-C_{ab}^{c}H^{(2)}_{ijc})q^{i}q^{j}a^{a}a^{b}+\frac{1}{3!}(\rho^{i}_{a}\partial_{i}H^{(1)}_{ibc}-C_{ab}^{d}H^{(1)}_{idc})q^{i}a^{a}a^{b}a^{c}$
$\quad-\frac{1}{4!}(\rho^{i}_{a}\partial_{i}H^{(0)}_{bcd}-C_{ab}^{e}H^{(0)}_{ecd})a^{a}a^{b}a^{c}a^{d}$
$=-\frac{1}{4}(\rho^{k}_{a}\partial_{k}H^{(2)}_{ijb}-C_{ab}^{c}H^{(2)}_{ijc})q^{i}q^{j}a^{a}a^{b}+\frac{1}{3!}(\rho^{i}_{a}\partial_{k}H^{(1)}_{ibc}-C_{ab}^{d}H^{(1)}_{idc})q^{i}a^{a}a^{b}a^{c}$
$\quad-\frac{1}{4!}(\rho^{i}_{a}\partial_{k}H^{(0)}_{bcd}-C_{ab}^{e}H^{(0)}_{ecd})a^{a}a^{b}a^{c}a^{d}.$ (101)
@@
${}^{A}\mathrm{d}^{\nabla}H^{(1)}-\mathrm{d}H^{(0)}=0,$ (113)
@@
$S_{F}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[\frac{1}{2}H^{(2)}_{ija}(\boldsymbol{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{A}}^{a}+\frac{1}{2}H^{(1)}_{iab}(\boldsymbol{X}){\bm{Q}}^{i}{\bm{A}}^{a}{\bm{A}}^{b}+\frac{1}{3!}H^{(0)}_{abc}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right].$ (115)
@@
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\boldsymbol{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+{\bm{W}}_{a}D{\bm{A}}^{a}+{\bm{Z}}_{i}^{\nabla}{\bm{Q}}^{i}+\rho^{i}_{a}(\boldsymbol{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}\right$.
$\left.+\frac{1}{3!}H^{(3)}_{ijk}(\boldsymbol{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{Q}}^{k}-\frac{1}{2}T_{ab}^{c}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{W}}_{c}\right$.
$\left.+\frac{1}{2}H^{(2)}_{ija}(\boldsymbol{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{A}}^{a}+\frac{1}{2}H^{(1)}_{iab}(\boldsymbol{X}){\bm{Q}}^{i}{\bm{A}}^{a}{\bm{A}}^{b}+\frac{1}{3!}H^{(0)}_{abc}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right].$ (116)